# 3DRealCar: An In-the-wild RGB-D Car Dataset with 360-degree Views

### v1

@@ 5.1 Experimental Setup
2D Car Parsing. In this task, we utilize DeepLabV3[11], DDRNet [23], SegFormer [71], and PointRend [33] to benchmark our dataset. To be specific, we split 80% of our car parsing maps in 3DRealCar as the training set and the rest of 20% as the testing set.
3D Reconstruction. In this task, we randomly choose 100 instances from each lighting condition in our dataset and split 80% of the views per instance as the training set and the rest of 20% as the testing set. Specifically, we employ Instant-NGP [48], 3DGS [31], GaussianShader [27], and 2DGS [24] to benchmark our dataset.
Novel View Synthesis. We finetune Zero123-XL [39] on our 3DRealCar dataset to enhance its generalization to real cars. Note that since the training of diffusion-based models needs entire objects centered on images, we use the images rendered by our trained 3D models as training images.
3D Generation. In this task, we exploit Dreamcraft3D [60] as our baseline. Dreamcract3D exploits Stable-Zero123 [2] as a prior source for providing 3D generative prior. By fine-tuning Stable-Zero123 on our dataset, we enable it to obtain car-specific prior so it generalizes well to real cars.
Table 4: Detection improvements by simulated data for corner-case scenes. We report the metric by calculating mAP@0.5 on CODA dataset [35].
Method Simulated Data
500 1000 2000 3000 4000 5000
YOLOv8n 0.261 0.299 0.312 0.357 0.386 0.386
YOLOv8s 0.336 0.371 0.366 0.403 0.413 0.435
YOLOv8m 0.347 0.354 0.362 0.416 0.419 0.441
YOLOv8l 0.358 0.375 0.381 0.411 0.413 0.458
@@
Corner-case 2D Detection. As shown in Table 4, we employ four variants of YOLOv8 serial models as our detectors for training. To evaluate the performance of models in corner-case scenes, we leverage the test part of the CODA dataset [35] as our testing set. In particular, when we increase the training simulated data from 500 to 5,000, the performance of detectors is also improved by a large margin. This phenomenon demonstrates that our simulated data is effective in improving a detector robust to corner-case scenes. We provide the visualizations of simulated corner-case scenes in Figure 6. The detailed simulation process and more visualizations can be seen in the supplementary.
@@
2D Car Parsing. We conduct benchmarks for car parsing maps of our dataset using recent image segmentation methods, such as DeepLabV3[11], PointRend[33], DDRNet[23], and SegFormer[71]. The quantitative performance for these methods on our dataset is summarized in Table 3. Visual comparisons are provided in Figure 7. Our high-quality dataset enables these methods to achieve promising performance, highlighting its potential for application in self-driving systems.
5.3 3D Tasks
3D Reconstruction. As depicted in Table 5, we show benchmark results of recent state-of-the-art 3D reconstruction methods, such as Instant-NGP [48], 3DGS [31], GaussianShader [27], and 2DGS [24] on our dataset. To the standard lighting condition, we can find that recent methods are capable of achieving PSNR more than 27 dB, which means these methods can reconstruct relatively high-quality 3D cars from our dataset. However, the reflective and dark condition results are lower than the standard. These two parts of our 3DRealCar bring two challenges to recent 3D methods. The first challenge is the reconstruction of specular highlights. Due to the particular property of cars, materials of car surfaces are generally glossy, which means it would produce plenty of specular highlights if cars are exposed to the sun or strong light. The second challenge is the reconstruction in a dark environment. The training images captured in the dark environment lose plenty of details for reconstruction. Therefore, how to achieve high-quality reconstruction results from these two extremely lighting conditions is a challenge to recent methods. 3D visualizations can be found on our project page.
Novel View Synthesis. As illustrated in the left part of Figure 8, we show visual comparisons of Zero123-XL [39] and our improved version by training on our dataset. As we can see, given input images, we use Zero123-XL and our improved version to synthesize novel views. In this figure, we can find that Zero123-XL prefers to generate synthetic results with unrealistic texture and geometry, due to the lack of prior for real objects. In contrast, our improved version of Zero123-XL can generate photorealistic geometry and texture, which demonstrates the effectiveness of our dataset.
3D Generation. As depicted in the right part of Figure 8, we visualize 3D generation results of the recent state-of-the-art single-image-to-3D method, Dreamcraft3D [60], along with its improved version by our dataset. This figure shows that Dreamcraft3D sometimes fails to generate complete geometry or realistic texture, due to the scarcity of the real car prior. As shown in Table 3, we also show quantitative comparisons of Dreamcraft3D and its improved version. CLIP-I means the similarity of rendered images with the original input. The quantitative and qualitative results indicate our dataset significantly improves 3D generation performance.
@@
2DGS [24] 27.58 0.9361 0.1084 23.19 0.8509 0.2041 22.63 0.9148 0.1681
Figure 8: Visualizations of novel view synthesis (left) and image-to-3D generation (right). we compare the results of the recent state-of-the-art method, Zero123-XL [39], Dreamcraft3D [60], and their improvement by training on our dataset.
@@ 6 Conclusion
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes]
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] We ensure our dataset would not bring negative societal impacts.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [No]
    2. (b)
      Did you include complete proofs of all theoretical results? [No]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We provide a URL in the abstract for researchers to download our dataset with detailed instructions.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Training details are described in Section 4. More details can be found in supplementary materials.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [No] They are public and we only use them for academic study.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [No]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Our dataset is collected under the consent of owners.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [No]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [No] Our dataset does not contain any potential participant risks.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [No] We collect our dataset by ourselves. We do not need to pay for participants.

### v2

@@ 5.1 Experimental Setup
2D Car Parsing. In this task, we utilize DDRNet [32], SegFormer [49], VMamba [25], and InternImage [47]to benchmark our dataset. To be specific, we split 80% of our car parsing maps in 3DRealCar as the training set and the rest of 20% as the testing set.
Neural Field-based Novel View Synthesis. In this task, we randomly choose 100 instances from each lighting condition in our dataset and split 80% of the views per instance as the training set and the rest of 20% as the testing set. Specifically, we employ recent state-of-the-art neural field methods, including Instant-NGP [29], 3DGS [18], GaussianShader [15], and 2DGS [13] to benchmark our dataset.
Diffusion-based Novel View Synthesis. We finetune Zero123-XL [22] on our 3DRealCar dataset to enhance its generalization to real cars. Note that since the training of diffusion-based models needs entire objects centered on images, we use the images rendered by our trained 3D models as training images.
Single Image to 3D Generation. In this task, we exploit Dreamcraft3D [40] as our baseline. Dreamcraft3D exploits Stable-Zero123 [39] as a prior source for providing 3D generative prior. By fine-tuning Stable-Zero123 on our dataset, we enable it to obtain car-specific prior so it generalizes well to real cars.
5.2 Evaluation Metrics
PSNR $\uparrow$: Peak Signal-to-Noise Ratio (PSNR) is a metric of the peak error between the original and a compressed or reconstructed image. Higher PSNR values indicate better image quality, with a higher similarity to original images.
SSIM $\uparrow$: Structural Similarity Index (SSIM)is a perceptual metric that considers changes in structural information, luminance, and contrast between the original and target image. Higher SSIM values indicate better performance.
LPIPS $\downarrow$: Learned Perceptual Image Patch Similarity (LPIPS) is a metric that uses deep learning models to assess the perceptual similarity between images. Lower LPIPS values indicate higher perceptual similarity. Unlike PSNR and SSIM, LPIPS leverages the capabilities of neural networks to better align with human visual perception.
mAP $\uparrow$: In the object detection task, mAP denotes mean Average Precision, a widely used metric to evaluate the performance of detection algorithms. It measures the accuracy of the detector by considering both the precision and recall at different thresholds. Higher mAP means better results.
@@ 5.2 Evaluation Metrics
Corner-case 2D Detection. To obtain a reliable detector for corner-case scenes, we simulate corner-case scenes with reconstructed cars and synthesize them within a background. Specifically, we leverage a recent large-scale self-driving dataset, Nuscenes [5] to provide background information. After obtaining a simulated corner-case dataset, we can use this dataset to train and obtain a reliable detector robust for corner-case scenes. As shown in Table 2, we employ YOLOv5 and YOLOv8 serial models, CO-DETR [52], and YOLOv12 [43] as our detectors for evaluation. To evaluate the performance of models in corner-case scenes, we leverage the test part of the CODA dataset [21] as our testing set. In particular, when we increase the training simulated data from 500 to 5,000, the performance of detectors improves by a large margin. This phenomenon demonstrates that our simulated data is effective in improving detection performance to corner-case scenes. We provide the visualizations of simulated corner-case scenes in Figure 6. The detailed simulation process and more visualizations can be seen in the supplementary.
Table 2: Detection improvements by simulated data for corner-case scenes. We leverage lightweight YOLO serials models and recent state-of-the-art models for evaluation. We report the metric by calculating mAP@0.5 on the CODA dataset [21].
Simulated Data YOLOv5n YOLOv5s YOLOv8n YOLOv8s CO-DETR YOLOv12x
1000 0.285 0.341 0.299 0.371 0.465 0.412
2000 0.304 0.357 0.312 0.366 0.481 0.441
3000 0.345 0.389 0.357 0.403 0.517 0.489
4000 0.357 0.408 0.386 0.413 0.551 0.531
5000 0.361 0.426 0.386 0.435 0.582 0.565
Method SegFormer DDRNet VMamba InternImage
mIOU $\uparrow$ 0.541 0.606 0.610 0.671
mAcc $\uparrow$ 0.652 0.732 0.734 0.786
@@
2D Car Parsing. We conduct benchmarks for car parsing maps of our dataset using recent segmentation methods, such as DDRNet[32], SegFormer[49], VMamba [25], and InternImage [47]. The quantitative performance for these methods on our dataset is summarized in Table 3. Visual comparisons are provided in Figure 7. Our high-quality dataset enables these methods to achieve promising performance, highlighting its potential for application in self-driving systems. In particular, our car parsing annotations encourage self-driving systems to recognize different components of cars in practical scenarios for safer automatic decisions. We believe that our detailed car parsing annotations could significantly contribute to advancing self-driving tasks.
@@
2DGS [13] 27.34 0.9341 0.1095 23.19 0.8509 0.2041 22.63 0.9148 0.1681
Pixel-GS [51] 27.67 0.9391 0.0994 24.81 0.8659 0.1541 23.54 0.9174 0.1617
3DGS-MCMC [19] 27.63 0.9382 0.0986 24.92 0.8681 0.1621 23.63 0.9198 0.1622
Table 5: Quantitative comparisons of SOTA 3D Generation method, Dreamcraft3D [40] and its improved version by trained on our dataset. CD denotes Chamfer Distance.
Method CLIP-I $\uparrow$ Hausdorff $\downarrow$ CD $\downarrow$
Dreamcraft3D 0.812 1.572 0.587
+our dataset 0.847 1.364 0.371
5.4 3D Tasks
Neural Field-based Novel View Synthesis. As depicted in Table 4, we show benchmark results of recent state-of-the-art neural field methods, such as Instant-NGP [29], 3DGS [18], GaussianShader [15], 2DGS [13], Pixel-GS [51], and 3DGS-MCMC [19] on our dataset. To the standard lighting condition, we can find that recent methods are capable of achieving PSNR more than 27 dB, which means these methods can reconstruct relatively high-quality 3D cars from our dataset. However, the reflective and dark condition results are lower than the standard. These two parts of our 3DRealCar bring two challenges to recent 3D methods. The first challenge is the reconstruction of specular highlights. Due to the particular property of cars, materials of car surfaces are generally glossy, which means it would produce plenty of specular highlights if cars are exposed to the sun or strong light. The second challenge is the reconstruction in a dark environment. The training images captured in the dark environment lose plenty of details for reconstruction. Therefore, how to achieve high-quality reconstruction results from these two extremely lighting conditions is a challenge to recent methods. 3D visualizations can be found on our project page. We hope these results can encourage subsequent research for the 3D reconstruction in low-quality lighting conditions.
Figure 8: Visualizations of diffusion-based novel view synthesis. we compare the results of the recent state-of-the-art diffusion-based method, Zero123-XL [22] and its improvement by training on our dataset. Our dataset provides car-specific prior for the generative model to generate more photorealistic car images.
Diffusion-based Novel View Synthesis. As illustrated in Figure 8, we show visual comparisons of Zero123-XL [22] and our improved version by training on our dataset. As we can see, given input images, we use Zero123-XL and our improved version to synthesize novel views. We can find that Zero123-XL prefers to generate unrealistic texture and geometry, due to the lack of prior for real objects. In contrast, our improved version of Zero123-XL can generate photorealistic geometry and texture, which demonstrates the effectiveness of our dataset.
Figure 9: Visualizations of single-image-to-3D generation. we compare the results of the recent state-of-the-art single-image-to-3D method, Dreamcraft3D [41] and is enhanced version by training on our dataset.
Single Image to 3D Generation. Not only do we enhance novel view synthesis for diffusion-based models, but we also demonstrate improvements in 3D generation. As depicted in Figure 9, we visualize 3D generation results of the recent state-of-the-art single-image-to-3D method, Dreamcraft3D [41], along with its improved version by our dataset. This figure shows that Dreamcraft3D sometimes fails to generate complete geometry or realistic texture, due to the scarcity of the real car prior. As shown in Table 5, we also show quantitative comparisons of Dreamcraft3D and its improved version. CLIP-I means the similarity of rendered images with the original input. The quantitative and qualitative results indicate our dataset significantly improves 3D generation performance typically in terms of geometry and texture. These results underscore the effectiveness of our 3DRealCar dataset.